A carousel has a radius of 1.65 m and a moment of inertia of 126 kg · m2 . A girl of mass 48.0 kg is standing at the edge of the carousel, which is rotating with an angular speed of 3.10 rad/s. Now the girl walks toward the center of the carousel and stops at a certain distance from the center d. The angular speed of the carousel is now 5.0 rad/s. How far from the center did the girl stop? __m
Answer #1

When standing at the edge, moment of inertia of system,
Ii = 126 + m*r^2
     = 126 + 48*1.65^2
      = 256.7 Kgm^2

final inertia,
If = 126 + m*d^2
     = 126 + 48*d^2

Use conservation of angular momentum:
Ii*wi = If*wf
256.7*3.1 = (126 + 48*d^2) * 5
795.8 = 630 + 240*d^2
d = 0.83 m
Answer : 0.83 m